In order to faster convert Account Receivable into cash, a company can sell its accounts receivable to another company, called a factor, at a discount.

Assume your average credit sales are $50 million per year and average accounts receivable are $5 million. A factoring company offers you an agreement that it will be buying your accounts receivable at a 2.20% discount. Your bank offers you a credit at R% per year (actual yearly rate).

What would be the minimum R% for which you would choose factoring? That is, what is the rate R% offered by the bank at or below which you would rather borrow money from the bank instead of factoring?

You should answer as the percentage. For example if the answer is 14.5%, then enter 14.5 as the answer.

(Hint: Refer to the Factoring Practice Problem. Assume 360 days in a year. Monthly bank interest rate to be paid is annual rate/12)

Answer:
Answer #1
By factoring , cash that will be immediately available is 5-(5*2.2%)=
4.89 mlns
& the cost of that funds made available= 5*2.2%=
0.11 mlns.
So, the
Average Cost %= 0.11/4.89=
2.25%
So, annual cost%=(1+0.0225)^12-1= 30.60%
So,if the
the rate R% offered by the bank is at or below 30.60% p.a. You would rather borrow money from the bank instead of factoring
as the average cost of factoring, in a year is 2.25%--that works out to 30.6% p.a.